OpenStudy (anonymous):

452% as a decimal and fraction in lowest form Also 0.1% and 0.75% 452% as a decimal and fraction in lowest form Also 0.1% and 0.75% @Mathematics

OpenStudy (karatechopper):

4.52, 4 13/25

OpenStudy (karatechopper):

0.001 1/1000

OpenStudy (karatechopper):

0.0075, 3/400

OpenStudy (anonymous):

4.52 4 13/25 o.1% = .001 = 1/1000 0.75% = .0075 = 75/10,000 = 3/400